Astronomy Report – February 20th

Crew 222 Astronomy Report 20Feb2020

Astronomy Report

Name: Eishi Kim

Crew: 222

Date: 2/20/2020


Robotic Telescope Requested (choose one) MDRS-WF

Objects to be Imaged this Evening: NGC 281, M42 and the Markarian’s chain are to be imaged as part of the astrophotography project.

Images submitted with this report: None

Problems Encountered: No calibration frames for M101 have been taken. Will follow Peter’s advice of using tonight’s calibration frames instead.


Solar Features Observed: The observatory was used today for a familiarization session in solar imaging. No significant phenomenon is to be reported.

Images submitted with this report: None

Problems Encountered: None

Sol Summary – February 20th

Crew 222 Sol Summary Report 20-02-2020


Summary Title: Routine day on Mars

Author’s name: Benjamin Auzou

Mission Status: Work in progress

Sol Activity Summary:

In the morning, the crew woke up with matinal workout session.

The purpose of today’s EVA was to find a path up to the Hab Ridge from the Kissing Camels. This was a success and the view from up here was splendid.

In the afternoon we worked on our own experiments.

In the end of the afternoon our HSO Nicolas organized a cooperation game for us.

Look Ahead Plan:

Tomorrow, we will explore the northern area of the Hab, we will make sure to pay more attention to the way we communicate during EVAs. Also with the weather report, we need to prepare the possibility of a rainy day on Saturday. For that we chose to send two of us on an EVA in the afternoon to bring back the LOAC, which can’t stay two days without a battery replacement.

Anomalies in work: none

Weather: mostly clear

Crew Physical Status: all members safe and sound

EVA: experiments maintenance and exploration of the Kissing Camels and Hab Ridge

Reports to be filed: Journalist report, GreenHab Report, Astronomy Report, EVA Report, EVA Request, Photos Report, Photo of the Day, Operations Report

Support Requested: Nothing to report

Operations Report – February 20th

Crew 222 Operations Report 20Feb2020

SOL: 4

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Wednesday generator:

Off at 6:30am

On at 9:00pm SOC was 83%

Thursday generator:

Off at 5:30am

On at 9pm SOC was XX%

Diesel Reading – 75%

Station Propane Reading – 34% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149356.82 units

Water (static tank) – 387 gallons

Static to Loft Pump used – yes

Water in GreenHab – 108.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 144.3

Beginning charge: 100%

Ending charge: 94%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: no

Hours: 132.0

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes

Opportunity rover used: no

Hours: 69.9

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes

Curiosity rover used: yes

Hours: 138.2

Beginning charge: 100%

Ending charge: 69%

Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes

Reason for use: Run during the engineer EVA for the battery.

Oil Added? no

ATV Fuel Used: N/A

Hours the ATVs were Used today: 10 minutes

Notes on ATVs: The second one from the Hab doesn’t have anymore battery.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com. All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the EVA. The spacesuit number 10 has stopped ventilating during the EVA this morning but the lights were still on. I tried to fix this during this afternoon and I’ll see if it works tomorrow.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

GreenHab Report – February 20th

Crew 222 GreenHab Report – 20Feb2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 20,7°C

Hours of supplemental light: NA

Daily water usage for crops: 9 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 108,5 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative: Today, I started by watering plants and then I put seeds in the gelatin i prepared yesterday, but it was too hot in the greenhab so i had to put them in the science dome, thos plants will be compared to thoses that will

grow in the culture tower moss. Then I prepared gelatin for the experiment on plants, called music for plants that i’ll start tomorrow. Tonight i watered plants again and harvested some basil and chives to put on our tonight pizza.

Harvest: 6g basil, 5g chiver

Support/supplies needed:

Journalist Report – February 20th

Crew 222 Journalist Report 20Feb2020

Author: Marie Bochard, Crew Journalist

Sol 4

Title: Precious water

The exploration continues! This morning, the EVA group headed down South to the Kissing Camels, a set of rocks between which the road passes through. Once there, we safely climbed over to the ridge, the hill that surpasses the road. The view up there was beautiful, on one side there was a vast plain with colored hilltops, on the other side there was a snowy mountain in the distance, between the high rock hills.

This afternoon, our crew Scientist, Nina, finished assembling her AeroSpring tower, the vertical culture tower. She also started to cultivate the plants that will grow in this tower, in a special moss that allows faster growth. We are hoping to see some sprouts before the end of our mission, and I will keep you posted!

Our Commander, Benjamin, also started putting together his experiment: the water-consumption control system. Using a camera linked to buttons, once someone wants to use water they just have to press a button before and the device will take a picture of the water-level before and after opening the tap. This monitoring is very important to our mission. Water is a rare resource here, and it has to be used with parsimony. Our biggest source of consumption is drinking: each crew member drinks up to 3L per day, since it is very dry here we are always thirsty. We also use water for cooking, as all our food comes dehydrated and has to be submerged in water before being cooked. The rest of the consumption comes from the dishes, hygiene and the toilet flush. We manage to use only 15 gallons of water per day in the Hab! Of course, in the GreenHab water is used to water the plants, and this system doesn’t take it into account since it is already monitored there.

Once again, Nicolas had prepared a relaxation session, which ended the afternoon well. Tonight is a special night: it’s pizza night! While writing our reports, we can all smell the oven heating up for our long-awaited dinner. Goodnight from every member of this amazing crew!

EVA Report – February 20th

Crew 222 EVA Report 20Feb2020

EVA # 4

Author: Marie BOCHARD (Journalist)

Purpose of EVA: Change batteries of LOAC and MegaARES, and then second exploration for the crew 222

Start time: 9:10 AM

End time: 11:21 AM

Narrative: First, we checked all the rovers and the water level.

We then took the rovers to the experiments site and changed the batteries for the LOAC and the MegaARES in teams of two. Once that was all done, we went back to the rovers with the used batteries and headed out to the Kissing Camels at 9:45.

We arrived soon after, and left the rovers on the side of the road before the Kissing Camels to go exploring. We walked towards the Camels looking for an old road going west towards the ridge. We explored the Kissing Camels for 15 minutes, and then we went back to the road we had found. We followed the road as best as we could until we reached the bottom of the Ridge. One crew member’s helmet was a little foggy, so we took breaks every 5 minutes to let his helmet defog. The hills in front of us were beautiful, and we looked for a way to climb up safely to the Ridge. We followed a dry stream of water until it was too steep, and continued walking up wherever the path was walkable for everyone.

Once we were up the Ridge, we admired the different views that were offered to us: the view of the plain up the Ridge, the view of the snowy mountain in the distance, and the view of the plain we came from and where we could see the rovers. We stayed up the Ridge for about 30 minutes, I made sure to have enough time to walk back down safely because climbing down is always more difficult than climbing up.

We then walked back to the rovers and headed back to the Hab. We entered the station at 11:21 AM.

Destination: Kissing Camels and Ridge

Coordinates (use UTM NAD27 CONUS): 518500-4249500

Participants: Marie BOCHARD (Journalist, EVA leader), Benjamin AUZOU (Commander), Eishi KIM (AST), Valentin PRUDHOMME (Crew engineer)

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Rovers : Curiosity & Deimos

Copyright © The Mars Society. All rights reserved. | Main Site